Question 105 pts

Holmes wanted to investigate whether deductive reasoning skills (x) predicted crime solving (y). He conducted a regression and found values α = 25 and b = 5. Lestrade scored a -5 on X. What would his Y value be predicted to be?

0

-5

5

-25
Answer #1

Solution :

b = 5

= 25

y = bx +

x = -5

Then y is,

y = 5 * (-5) + 25 = -25 + 25 = 0

Y value be predicted to be 0